# FalconFS: Distributed File System for Large-Scale Deep Learning Pipeline

### v1

@@ Abstract
Client-side metadata caching has long been considered an effective method for accelerating metadata operations in distributed file systems (DFSs). However, we have found that client-side state (e.g., caching) is not only ineffective but also consumes valuable memory resources in the deep learning pipelines. We thus propose FalconFS, a DFS optimized for deep learning pipelines with the stateless-client architecture. Specifically, instead of performing client-side path resolution and caching, FalconFS efficiently resolves paths on the server side using hybrid metadata indexing and lazy namespace replication. FalconFS also boosts server concurrency with concurrent request merging and provides easy deployment with VFS shortcut. Evaluations against CephFS and Lustre show that FalconFS achieves up to 5.72$\times$ throughput for small file read/write and up to 12.81$\times$ throughput for deep learning model training. FalconFS has been running in Huawei autonomous driving system’s production environment with 10,000 NPUs for one year.
@@ 1 Introduction
However, the POSIX interface and tree-structured directory organization of DFSs are a double-edged sword, particularly in the context of deep learning (DL) workloads. On one hand, the POSIX interface is valued for its generality and convenience, facilitating integration with existing applications and frameworks. On the other hand, the POSIX interface and hierarchical directory organization are not well-suited for distributed environments, leading to inefficiencies in many key use cases. Specifically, the tree-structured organization requires frequent path resolution in DFSs. Before any file operation, the DFS client must resolve the complete path to locate the target file’s inode, which involves verifying the existence and permissions of every directory along the path (Figure 1). As modern DFSs typically distribute directory metadata across multiple metadata servers for scalability, path resolution entails multiple round-trips between the client and metadata servers. This leads to significant request amplification: a single file operation entails multiple network requests.
@@
Figure 2: CephFS performance of randomly traversing 64 KiB files in a large directory tree under different client metadata cache sizes. The left y-axis represents the read throughput and the right y-axis represents the number of requests sent to the MDSs (i.e., lookup and close). The dashed line denotes the file number.
@@
When accessing such massive directory structures, the stateful-client architecture faces an inherent trade-off: either consume a huge amount of client memory to cache the large directory tree or suffer a severe performance drop due to request amplification. To present this trade-off, we evaluate how the metadata cache size affects the performance of random file traversal in a large directory tree, as shown in Figure 2 (further explained in §2.3). Compared to a cache that stores all directories, a cache only 10% as large results in a 32.5% throughput reduction, primarily because the lookup requests increase by 1.50$\times$. The dilemma is further exacerbated by the fact that even caching 10% of the directories is prohibitively expensive in production, considering the large scale directory tree and the client number. Even when 90% of directories are cached, an average of 1.70 network hops (i.e., the lookup requests in the figure) are required to complete a file open.
In this paper, we propose the stateless-client architecture, which achieves one-hop access for most file operations in DL workloads while requiring no client-side caching. The core difference is that stateless client shifts path resolution to the server side. However, to realize the stateless-client architecture for an efficient DFS, two key problems remain to be addressed. First, to enable one-hop access for most file operations, the client needs an approach to finding the correct metadata server that stores the target file’s inode. Second, as the path resolution is now on the server side, each metadata server should be able to resolve path locally, without involving additional network hop when processing client’s file operation request.
@@
To address the second problem of resolving path locally, FalconFS proposes lazy namespace replication, which replicates the filesystem’s whole namespace (i.e., the directory tree) to all metadata servers (§4.3). To reduce maintenance overhead, modifications to the namespace are lazily synchronized and an invalidation-based mechanism is adopted for concurrency control.
@@
Evaluations show that FalconFS completely eliminates request amplification and improves performance of large directory tree random traversal by up to 4.72$\times$ over CephFS [48] and up to 3.34$\times$ over Lustre [4]. End-to-end experiments show that FalconFS improves the throughput of DL model training by up to 11.81$\times$ and 1.23$\times$ over CephFS and Lustre, respectively. FalconFS has been deployed in Huawei’s AI clusters with 10,000 NPUs for data labeling and model training of the autonomous driving solution for one year.
@@
  • Building FalconFS, a high-performance DFS designed for DL workloads and a thorough evaluation on it.
@@ 2.1 Deep Learning Pipeline
Figure 3 illustrates the architecture of the deep learning (DL) training pipeline employed in Huawei’s autonomous driving (AD) systems. The pipeline consists of four key stages: The ingestion stage collects raw data from real-world environments. The labeling stage generates labels for the raw data with a sequence of model inference tasks, including moving objects detection, lane detection, traffic sign detection, etc. Then, the training stage uses a labeled dataset to train the target model for vehicle deployment. In the final archiving stage, the dataset is moved into low-cost storage systems such as cloud data lakes for future reference [49, 30, 51, 16]. Similar architectures also exist in other DL workloads [51, 1, 31].
@@ 2.2 Workload Patterns in DL pipelines
We take the DL pipeline for autonomous driving as an example to present unique workload patterns in DL pipelines.
@@
The autonomous driving pipeline consumes multimodal data, including images, point clouds, etc. During labeling and training, these objects are stored as individual files, whose size ranges from a few KiB to a few MiB, mostly within 256KiB. In production, an inflight dataset scales up to hundreds of petabytes and is composed of over 300 billion small files. These files are grouped into directories by timestamps, vehicle ID, camera ID, etc., forming a directory tree with billions of directories and large directory sizes. The huge number of files and directories stresses the DFS’s metadata scalability and performance.
@@
In the labeling stage, the inference tasks read and write files from/to FalconFS in a pipeline, involving massive small file IO and directory lists. To fully utilize GPU’s parallelism, data objects are accessed and processed in batches, resulting in burst file access in the same directory, which can lead to instantaneous load imbalance on the metadata servers and downgrade the performance, which we further discuss in §2.4.
@@ 2.3 Challenge 1: Lookup Tax
To illustrate this dilemma, we replay a trace of Resnet-50 model training on a small dataset, comprising 10 million 64KiB files in 1 million 7-level directories, stored in a CephFS instance with 4 MDSs and 12 OSDs A total of 512 IO threads iterate through the files in random order. As shown in Figure 2, the client cache size significantly impacts both read throughput and the number of remote requests generated.
The experiment presents that the read throughput is highly sensitive to the client’s metadata cache size. Compared to caching 10% directories, caching all directories achieves 1.46$\times$ higher throughput. Furthermore, throughput increases significantly as the cache size grows from 10% to 100%, indicating that optimal performance requires allocating sufficient memory to cache all directories. Otherwise, performance degrades proportionally with reductions in cache size.
@@
The DL training workload’s sensitivity to cache size comes from its random access nature. During each epoch, training tasks access every data object exactly once in a random order. When all directories are cached, all directory lookups are served by the cache, reducing each file open operation to a single metadata lookup request. However, with a small cache size, the LRU policy preferentially retains near-root directories, while the hit rate of last-level directories — which constitute 90% of accesses in the experiment —- is proportional to the cache size. As shown in Figure 2, smaller caches correlate with increased lookup requests due to cache misses. This effect causes request amplification: each file open operation triggers multiple lookup requests to the metadata server, bottlenecking read throughput.
@@
In production environments, caching a significant portion of working-set directories proves prohibitively expensive. For instance, a production cluster in Huawei can scale over 1000 client nodes and a typical production dataset contains billions of directories. Given that in Linux VFS, caching a directory takes 800 bytes (608 bytes for inode and 192 bytes for dentry), caching 10% of 1 billion directories on each node would require 80 GiB per node and 80 TiB in total, which is prohibitively expensive.
2.4 Challenge 2: Burst IO
@@
During the labeling stage, inference tasks load and store data objects to working directories using batched file operations. This batched IO pattern causes instantaneous load imbalances, which limit performance scalability. We illustrate this issue by performing burst file access on a CephFS cluster with 4 MDSs and 12 OSDs and present the results in Figure 4.
@@
We observe that when the burst size exceeds the IO parallelism of the tasks, the DFS’s performance degrades. Analysis of load distribution (Figure 4) shows that at a burst size of 100, the MDSs experience severe load imbalance during read operations. This issue arises because CephFS tends to store metadata for files within the same directory together; consequently, burst operations on files in the same directory can congest a single MDS, leading to performance bottlenecks. Similar patterns also exist in other DFSs like [15, 36, 27, 46, 34].
@@
Figure 4: CephFS performance of accessing 64KiB files with different burst size. Large burst size leads to load imbalance and performance degradation. Figure 4 shows the load variance of the MDSs when performing read operations with burst size 100.
@@ 3 Proposal: DFS with Stateless Client
Given the challenges brought by DL workloads, we propose the stateless-client architecture for DFS under DL workloads. The stateless-client architecture abandons the cache on the client side and moves path resolution to the server side. The proposal is based on the following three observations. (1) The server has more memory resources than the clients due to the tight resource budget of the computing node as we described in §2.2. (2) A typical DL cluster usually has far more clients than metadata servers. In Huawei’s deployment environment, the ratio of client machines to server machines exceeds 40:1, and each client machine can host multiple clients. Considering that all clients share the same dataset in DL pipelines, a server-side dentry can serve more clients than a client-side cache. (3) Stateful-client DFSs typically use VFS dcache and inode cache to cache directory attributes [48, 4, 3, 15], which takes 800 bytes for each directory. On the server side, a directory entry can be maintained in a custom format which takes less than 100 bytes. However, to realize the stateless-client architecture for an efficient DFS, two key problems remain.
@@
However, none of these indexing approaches meets of requirements of an efficient DFS with the stateless-client architecture, including no client cache, no extra network hops, and practical support of directory renames. Nevertheless, these approaches inspire us to build a hybrid indexing of hashing and path-walking to solve the problem.
@@
Second, as the path resolution is now on the server side, each metadata server should be able to resolve the path locally, without involving another network hop when processing the client’s file operation request. Since DFS scatters directory entries across multiple metadata servers for scaling out, the dentry information of components in a path is also scattered on multiple servers. A server resolving the path thus needs to communicate with other servers to complete the resolution, which can be expensive.
@@
With hybrid metadata indexing and lazy namespace replication, we build FalconFS with the stateless-client architecture, which eliminates client-side caching while providing one-hop access for most operations in DL workloads.
@@ 4.1 FalconFS Overview
Figure 5 shows the FalconFS architecture, consisting of Client Module, MNode, Coordinator, and File Store.
@@
File Store is a distributed block storage storing file data. File blocks are indexed by hashing file ID combined with block offset.
@@
FalconFS replicates the file system directory structure across all MNodes, enabling each MNode to resolve file paths and check permissions locally. The namespace replica contains entries of directory attributes required for path resolution, i.e., dentries in Table 1, and does not include file attributes. With one billion directories, the storage footprint for the namespace replica is less than 100 GiB per MNode, which is acceptable.
@@
In contrast to directories, we distribute all the file metadata, i.e., inodes in Table 1, across the metadata servers by hybrid metadata indexing for scaling up metadata capacity and throughput.
@@ 4.2.1 Hybrid Indexing Methods
Hybrid metadata indexing adopts filename hashing in the common case. However, for generality, hybrid metadata indexing considers situations where filename hashing may bring uneven inode distribution, and adopts two fallback mechanisms to mitigate the problem. Figure 6 presents an overview.
@@
To accelerate the path resolution, FalconFS adopts filename hashing as the indexing methods for most cases. With filename hashing, all files are placed via the hashing of the filename. Such an indexing method is easy and efficient — it requires no client state and support efficient rename. However, filename hashing cannot guarantee that the files are distributed evenly among different servers, potentially causing load-imbalance issues.
Fortunately, we find that DL workloads’ large directory size facilitate the filename hashing, significantly reducing the occurrence of load imbalance. Specifically, DL datasets usually has large directory size, which can be multiple times larger than the number of MNodes. We analyze Huawei’s in-production datasets and popular open-sourced datasets that contains more than ten thousand files [17, 13, 11, 26, 44], finding their directory size ranging from several hundreds to hundreds of thousands. According to the law of large numbers, given such large directory sizes, the distribution of files in each directory to MNodes is likely to be uniform. Consequently, the entire namespace is statistically uniformly distributed as the superposition of the distributions for each directory. This forms the basis of using filename hashing as the indexing method for the common cases in DL workloads.
@@
  • Path-walk redirection. To place files with hot filenames, FalconFS calculates the hashing value by not only the filename, but also its parent directory ID. Thus, even if a hot filename exists in many directories, the corresponding files are placed on different MNodes. When clients identify a file marked for path-walk redirection in the exception table, they send requests to random MNodes. The receiving node utilizes its local namespace replica to walk and obtain the parent directory ID, and forwards the request to the target node determined by hashing both filename and parent directory ID.
  • Overriding redirection. When hash variance causes uneven filename distribution across nodes, FalconFS can reassign selected filenames to designated nodes, shifting load from overloaded to underutilized nodes. These placement overriding are maintained in the exception table. Clients encountering files marked for such redirection in the exception table send requests directly to their designated MNodes.
FalconFS maintains copies of the exception table on each MNode, clients and the coordinator. The coordinator updates the exception table according to the scheduling policy detailed in §4.2.2. Once the exception table is updated, the latest table is eagerly pushed to all MNodes, and clients lazily fetch the updates from MNodes when they get responses to operation requests. The lazy fetching mechanism creates a time window where clients may operate with stale exception tables and directing requests to incorrect nodes. However, MNodes validate all requests by checking its local exception table and forward misdirected requests to the proper destinations.
A constant number of exception table entries are sufficient to balance the distribution of inodes for arbitrary directory structures. We provide a theoretical analysis in §A.1.
4.2.2 Statistical load balancing.
The coordinator uses the statistics periodically reported by MNodes to make re-balancing decisions.

### v2

@@ Abstract
[…]
@@ 1 Introduction
However, the POSIX interface and tree-structured directory organization of DFSs are a double-edged sword, particularly in the context of deep learning (DL) workloads. On one hand, the POSIX interface is valued for its generality and convenience, facilitating integration with existing applications and frameworks. On the other hand, the POSIX interface and hierarchical directory organization are not well-suited for distributed environments, leading to inefficiencies in many key use cases. Specifically, the tree-structured organization requires frequent path resolution in DFSs. Before any file operation, the DFS client must resolve the complete path to locate the target file’s inode, which involves verifying the existence and permissions of every directory along the path (Fig. 1). As modern DFSs typically distribute directory metadata across multiple metadata servers for scalability, path resolution entails multiple round-trips between the client and metadata servers. This leads to significant request amplification: a single file operation entails multiple network requests.
@@
Figure 2: CephFS performance of randomly traversing 64 KiB files in a large directory tree under different client metadata cache sizes. The left y-axis represents the read throughput, and the right y-axis represents the number of requests sent to the MDSs (i.e., lookup and close). The dashed line denotes the file number.
@@
[…]
In this paper, we propose the stateless-client architecture, which achieves one-hop access for most file operations in DL workloads while requiring no client-side caching. The core difference is that the stateless client shifts path resolution to the server side. However, to realize the stateless-client architecture for an efficient DFS, two key problems remain to be addressed. First, to enable one-hop access for most file operations, the client needs an approach to finding the correct metadata server that stores the target file’s inode. Second, as the path resolution is now on the server side, each metadata server should be able to resolve paths locally, without involving additional network hops when processing clients’ file operation requests.
@@
To address the second problem of resolving path locally, FalconFS proposes lazy namespace replication, replicating the whole namespace (i.e., the directory tree) to all metadata servers (§4.3). To amortize synchronization overhead, namespace updates are lazily synchronized, and an invalidation-based mechanism is adopted for concurrency control.
@@
[…]
@@
  • Building FalconFS, a high-performance DFS designed for DL workloads, and a thorough evaluation on it.
@@ 2.1 Deep Learning Pipeline
Fig. 3 illustrates the architecture of the deep learning (DL) training pipeline employed in Huawei’s autonomous driving (AD) systems. The pipeline consists of four key stages: The ingestion stage collects raw data from real-world environments. The labeling stage generates labels for the raw data with a sequence of model inference tasks, including moving objects detection, lane detection, traffic sign detection, etc. Then, the training stage uses a labeled dataset to train the target model for vehicle deployment. In the final archiving stage, the dataset is moved into low-cost storage systems such as cloud data lakes for future reference [50, 31, 52, 16]. Similar architectures also exist in other DL workloads [52, 1, 32].
@@ 2.2 Workload Patterns in DL pipelines
We take the DL pipeline for autonomous driving in Huawei as an example to present its unique workload patterns.
@@
The autonomous driving pipeline consumes multimodal data, including images, point clouds, etc. During labeling and training, these objects are stored as individual files, whose size ranges from a few KiB to a few MiB, mostly within 256KiB. In production, an in-flight dataset scales up to hundreds of petabytes and is composed of over 300 billion small files. These files are grouped into directories by timestamps, vehicle ID, camera ID, etc., forming a directory tree with billions of directories and large directory sizes. The huge number of files and directories stresses the DFS’s metadata scalability and performance.
@@
In the labeling stage, the inference tasks read and write files from/to FalconFS in a pipeline, involving massive small file IO and directory lists. To fully utilize the GPU’s parallelism, data objects are accessed and processed in batches, resulting in burst file access in the same directory, which can lead to instantaneous load imbalance on the metadata servers and downgrade the performance, which we further discuss in §2.4.
@@ 2.3 Challenge 1: Lookup Tax
To illustrate this dilemma, we replay a trace of Resnet-50 model training on a small dataset, comprising 10 million 64KiB files in 1 million 7-level directories, stored in a CephFS instance with four MDSs and twelve OSDs [^2] Ceph MDS (metadata server daemon) and OSD (object storage daemon) can be considered as CephFS’s metadata server and data node, respectively.. A total of 512 IO threads iterate through the files in random order. As shown in Fig. 2, the client cache size significantly impacts both read throughput and the number of remote requests generated.
[…]
@@
The DL training workload’s sensitivity to cache size comes from its random access nature. During each epoch, training tasks access every data object exactly once in a random order. When all directories are cached, all directory lookups are served by the cache, reducing each file open operation to a single metadata lookup request. However, with a small cache size, the LRU policy preferentially retains near-root directories, while the hit rate of last-level directories — which constitute 90% of accesses in the experiment —- is proportional to the cache size. As shown in Fig. 2, smaller caches correlate with increased lookup requests due to cache misses. This effect causes request amplification: each file open operation triggers multiple lookup requests to the metadata server, bottlenecking read throughput.
@@
In production environments, caching a significant portion of working-set directories proves prohibitively expensive. For instance, a production cluster in Huawei can scale over 1000 client nodes, and a typical production dataset contains billions of directories. Given that in Linux VFS, caching a directory takes 800 bytes (608 bytes for inode and 192 bytes for dentry), caching 10% of 1 billion directories on each node would require 80 GiB per node and 80 TiB in total, which is prohibitively expensive.
2.4 Challenge 2: Transient Skewness
@@
During the labeling stage, inference tasks scan and load/store data objects in a per-directory manner, accessing all the files in one directory and then another. This per-directory IO pattern causes transient load imbalances, which limit performance scalability. We illustrate this issue by performing per-directory file access on a CephFS cluster with four MDSs and twelve OSDs and present the results in Fig. 4.
@@
We observe that when the directory size exceeds the IO parallelism of the tasks, the DFS’s performance degrades. Analysis of load distribution (Fig. 4) shows that at a directory size of 100, the MDSs experience severe load imbalance during read operations. This issue arises because CephFS tends to store metadata for files within the same directory together; consequently, burst operations on files in the same directory can congest a single MDS, leading to performance bottlenecks. Similar patterns also exist in other DFSs like [15, 37, 28, 47, 35].
@@
Figure 4: CephFS performance of accessing 64KiB files with different burst size. Large burst size leads to load imbalance and performance degradation. Fig. 4 shows the load variance of the MDSs when performing read operations with burst size 100.
@@ 3 Proposal: DFS with Stateless Client
Given the challenges brought by DL workloads, we propose the stateless-client architecture for DFS under DL workloads. The stateless-client architecture abandons the cache on the client side and moves path resolution to the server side. The proposal is based on the following three observations. (1) A typical DL cluster usually has far more clients than metadata servers. In Huawei’s deployment environment, the ratio of clients to servers exceeds 40:1. Considering that all clients share the same dataset in DL pipelines, a server-side dentry can serve more clients than a client-side cache. (2) Stateful-client DFSs typically use VFS dcache and inode cache to cache directory attributes [49, 4, 3, 15], which takes 800 bytes for each directory. On the server side, a directory entry can be maintained in a custom format that takes less than 100 bytes. (3) The server has more memory resources than the clients due to the tight resource budget of the computing node, as we described in §2.2. However, to realize the stateless-client architecture for an efficient DFS, two key problems remain.
@@
However, none of these indexing approaches meet the requirements of an efficient DFS with the stateless-client architecture, including no client cache, no extra network hops, and practical support of directory renames. Nevertheless, these approaches inspire us to build a hybrid indexing of hashing and path-walking to solve the problem.
@@
Second, as the path resolution is now on the server side, each metadata server should be able to resolve the path locally, without involving another network hop when processing the client’s file operation request. Since DFS scatters directory entries across multiple metadata servers for scaling out, the dentry information of components in a path is also scattered across multiple servers. A server resolving the path thus needs to communicate with other servers to complete the resolution, which can be expensive.
@@
We build FalconFS with the stateless-client architecture, which eliminates client-side caching while providing one-hop access for most operations in DL workloads with hybrid metadata indexing and lazy namespace replication.
@@ 4.1 FalconFS Overview
Fig. 5 shows the FalconFS architecture, consisting of Client Module, MNode, Coordinator, and File Store.
@@
File Store is a distributed block storage system that stores file data. File chunks are distributed across a set of file store nodes that use local file systems on SSDs to store data.
@@
FalconFS replicates the file system directory structure across all MNodes, enabling each MNode to resolve file paths and check permissions locally. The namespace replica contains entries of directory attributes required for path resolution, i.e., dentries in Tab. 1, and does not include file attributes. With one billion directories, the storage footprint for the namespace replica is less than 100 GiB per MNode, which is acceptable.
@@
In contrast to directories, we distribute all the file metadata, i.e., inodes in Tab. 1, across the metadata servers by hybrid metadata indexing for scaling up metadata capacity and throughput.
@@ 4.2.1 Hybrid Indexing Methods
Hybrid metadata indexing adopts filename hashing in the common case. However, for generality, hybrid metadata indexing considers situations where filename hashing may bring uneven inode distribution, and adopts two fallback mechanisms to mitigate the problem. Fig. 6 presents an overview.
@@
To accelerate the path resolution, FalconFS adopts filename hashing as the indexing method for most cases. With filename hashing, all files are placed via the hashing of the filename. Such an indexing method is easy and efficient — it requires no client state and supports efficient rename. However, filename hashing cannot guarantee that the files are distributed evenly among different servers, potentially causing load-imbalance issues.
Fortunately, we find that DL workloads’ large directory size facilitates the filename hashing, significantly reducing the occurrence of load imbalance. Specifically, DL datasets usually have a large directory size, which can be multiple times larger than the number of MNodes. We analyze Huawei’s in-production datasets and popular open-sourced datasets that contain more than ten thousand files [17, 13, 11, 27, 45], finding their directory size ranging from several hundred to hundreds of thousands. According to the law of large numbers, given such large directory sizes, the distribution of files in each directory to MNodes is likely to be uniform. Consequently, the entire namespace is statistically uniformly distributed as the superposition of the distributions for each directory. This forms the basis of using filename hashing as the indexing method for the common cases in DL workloads.
@@
  • Path-walk redirection. To place files with hot filenames, FalconFS calculates the hashing value by not only the filename, but also its parent directory ID. Thus, even if a hot filename exists in many directories, the corresponding files are placed on different MNodes. When clients identify a file marked for path-walk redirection in the exception table, they send requests to random MNodes. The receiving node utilizes its local namespace replica to walk the path, obtains the parent directory ID, and forwards the request to the target node determined by hashing both the filename and the parent directory ID.
  • Overriding redirection. When hash variance causes uneven filename distribution across nodes, FalconFS can reassign selected filenames to designated nodes, shifting load from overloaded to underutilized nodes. These placement overrides are maintained in the exception table. Clients encountering files marked for such redirection in the exception table send requests directly to their designated MNodes.
FalconFS maintains copies of the exception table on each MNode, clients, and the coordinator. The coordinator updates the exception table according to the scheduling policy detailed in §4.2.2. Once the exception table is updated, the latest table is eagerly pushed to all MNodes, and clients lazily fetch the updates from MNodes when they get responses to operation requests. The lazy fetching mechanism creates a time window where clients may operate with stale exception tables and direct requests to incorrect nodes. However, MNodes validate all requests by checking their local exception table and forward misdirected requests to the proper destinations.
A constant number of exception table entries is sufficient to balance the distribution of inodes for arbitrary directory structures. We provide a theoretical analysis in §A.1.
4.2.2 Statistical Load Balancing
The coordinator uses the statistics periodically reported by MNodes to make rebalancing decisions.